OpenStudy (anonymous):

Help (4a^2b)^3 a. 12a^6b^3 b. 64a^5b^3 c. 12a^5b^3 d. 64a^6b^3

OpenStudy (anonymous):

\[\Large (4a^2b)^3\] Correct?

OpenStudy (uri):

4^3? =4*4*4=64

OpenStudy (uri):

(a^2)^3=a^6

OpenStudy (acxbox22):

b^3=b^3

OpenStudy (acxbox22):

combine the 3

OpenStudy (uri):

So, what would be the answer? @MUNCH300

OpenStudy (anonymous):

Wouldnt It Be 12a^6b^3

OpenStudy (acxbox22):

yes

OpenStudy (uri):

@MUNCH300 :)